OpenStudy (anonymous):

write as a single fraction. simplify. -8b-3y/ 8b + 5b+2y/ 2b +1

OpenStudy (yolo_king):

Would it be written like this? \[-8b-3\frac{ y }{ 8b }+5b+2 \frac{ y }{ 2b}+1\]

OpenStudy (anonymous):

yes exactly =)

OpenStudy (yolo_king):

Ok C: then the answer would be \[-3b+\frac{ 5y }{ 8b }+1\]

OpenStudy (yolo_king):

^Simplest Form^

OpenStudy (anonymous):

@Luigi0210 sorry this post is easier to get to I didn't realize I had posted it again =o

OpenStudy (anonymous):

Thank you @YOLO_KING =)

OpenStudy (yolo_king):

You bet C: please medal mehhh qq

OpenStudy (anonymous):

it says it has to be written as a single fraction =O

OpenStudy (yolo_king):

:o Whaaaaaaa?

OpenStudy (yolo_king):

.-. thats not possible hold on

OpenStudy (luigi0210):

Well I'll post it on both then.. Actually the common factors would be 8b, so just multiply the second by 4 to get 8b, and the 1 will be 8b/8b, like so: \(\LARGE -\frac{8b-3y}{8b} +\frac{5b+2y}{2b} (4) + \frac{8b}{8b} \) Simplify: \(\LARGE -\frac{8b-3y}{8b} +\frac{20b+8y}{8b} + \frac{8b}{8b} \) Now get it all over a common denominator like so: \(\LARGE \frac{(20b-8b)+(8y+3y)+(8b)}{8b}\) I also distributed the negative out So it'll just be \(\LARGE \frac{20b+11y}{8b}\) Make sense?

OpenStudy (yolo_king):

^qq^ dang

OpenStudy (anonymous):

lol I gave medals to both because you both help so much you guys are the best =) thank you

OpenStudy (yolo_king):

Well I see your in good hands @Cassandra_Lea_96 see ya qq